# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 2. Geometric properties near the boundary
Furthermore, we have $|\nabla\phi_{i}|<\frac{1}{100}$, $|\nabla^{2}\phi_{i}|<\widetilde{C}(\Omega)$ for some $\widetilde{C}(\Omega)>\frac{4}{d}$, and for any $1\leq i\leq m_{1}$,
@@
This is a direct consequence of the fact that $|\nabla\phi_{i}|<\frac{1}{100}$ and the mean value inequality. Indeed, given $x=x_{i}^{0}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x_{i}^{0})$, if there exists $x^{\prime}=x_{i}^{0}+x^{\prime}_{1}e^{1}_{i}+x^{\prime}_{2}e^{2}_{i}-x^{\prime}_{3}n(x_{i}^{0})\in B(x,\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2})))$ such that $x^{\prime}\in\partial\Omega\cap B(x^{0}_{i},3d))$, we have by Lemma 2.1 that
@@
Remark 3.
Notice that by the triangle inequality we have $\Omega_{\frac{d}{8}}\subset\cup_{i=1}^{m_{1}}B(x_{i}^{0},\frac{d}{4})$, where $\Omega_{\epsilon}:=\{x\in\Omega|\ d(x,\partial\Omega)<\epsilon\}$.
Remark 4.
It is known that any bounded domain with $C^{2}$ boundary satisfies the uniform interior sphere condition. That is, there exists $d_{r}$ such that for any $x\in\partial\Omega$, one can find a ball with radius $d_{r}>0$ in $\Omega$ that intersects with $\partial\Omega$ only on $x$. In this article, we choose $d$ from Lemma 2.1 to be smaller than $d_{r}$ and take larger $\widetilde{C}(\Omega)$ such that $\widetilde{C}(\Omega)\geq\frac{1}{2d_{r}}$.
@@
By Lemma 2.1 and Remark 3, there exist $m_{1}$ and $\{x_{i}^{0}\}_{i=1}^{m_{1}}$ such that
@@
Clearly, we have $z^{\prime\prime}-z_{q_{1}}=W_{1,//}+W_{1,\perp}$, $W_{1,//}\mathbin{\!/\mkern-5.0mu/\!}Z_{2}\mathbin{\!/\mkern-5.0mu/\!}x-y$, and $W_{1,\perp}\perp Z_{2}$.
@@
Suppose that $\Omega$ satisfies Assumption A with $\delta$, $d$ as given in Lemma 2.1. Then, $\Omega-\Omega_{\frac{d}{8}}$ is path-wise connected, that is, given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$, there exists a continuous curve in $\Omega-\Omega_{\frac{d}{8}}$ connecting $x^{\prime}_{1}$ and $x^{\prime}_{2}$.
@@
Given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$. Since the domain $\Omega$ is connected, $\Omega$ is path-wise connected. Hence, there exist $l\in\mathbb{N},\{\widetilde{x}_{i}\}_{i=1}^{l}\in\Omega$ such that the following curve
@@
As Remark 3 shows, we have
@@
We also notice that by using the same argument as (2.11), we have $|n(x^{0}_{k(j)})-n(x^{0}_{k(j+1)})|<\frac{1}{50}$ for $1\leq j\leq l-1$, which we deduce that
@@
all the assumption of Lemma 2.3 is satisfied. Hence, we deduce that
@@
for $i=1,2$. Indeed, if $x^{\prime}_{i}\in\Omega_{\frac{d}{8}}$, then by Lemma 2.1, there exists $1\leq l(i)\leq m_{1}$ such that $x^{\prime}_{i}\in B(x_{l(i)}^{0},\frac{d}{4})$ and we can parameterize the boundary of $\Omega$ near $x^{0}_{l(i)}$ as below:
@@
By mean value theorem, there exist $\zeta^{\prime}$ such that
@@
(2.20) $\begin{split}&\mathbb{O}(w_{1}):=\{y\in\mathbb{Y}\mid\mathrm{There\ exists\ a\ zigzag}\,\ \{w_{1},u_{1},...,u_{s},y\}\ni\,u_{i}\in\mathbb{Y}\,\forall\,i=1,...,s,\\ &d(\overline{w_{1}u_{1}},\partial\Omega)>\frac{d}{10},d(\overline{u_{i}u_{i+1}},\partial\Omega)>\frac{d}{10}\forall\,1\leq i\leq s-1,d(\overline{u_{s}y},\partial\Omega)\>\frac{d}{10}\}.\end{split}$
@@
$\overline{\xi}:=\sup\{0\leq s\leq 1\mid d(\mathbb{O}(w_{1}),\overline{\phi}(s))\leq\frac{d}{15}\}.$
@@
Thus, we have that $w_{4}\in\mathbb{O}(w_{1})$. This implies that $\frac{d}{16}\leq d(\mathbb{O}(w_{1}),w_{3})\leq d(w_{4},w_{3})<\frac{d}{16}$, which is a contradiction. We conclude that $\Omega-\Omega_{\frac{d}{8}}\subset B\left(\mathbb{O}(w_{1}),\frac{d}{16}\right)$. Therefore, $\mathbb{O}(w_{1})$ contains $\mathbb{Y}$, and $w_{2}\in\mathbb{O}(w_{1})$, which shows that there exists a zigzag between $w_{1}$ and $w_{2}$ satisfying the conditions mentioned in Lemma 2.5.
@@
We invoke the Remark 4 and use the uniform interior sphere condition to deduce that for any $0\leq s\leq\frac{\sin{\theta_{i}}}{\widetilde{C}(\Omega)}$, we have
@@
For the (2.30), we again use the uniform interior sphere condition to deduce that for any $y\in\partial\Omega$ and $w\in\mathbb{R}^{3}$ with $n(y)\cdot w<0$, $\theta_{w}:=\frac{\pi}{2}-\arccos{|n(y)\cdot w|}$, we have
@@ 3. Operator estimates and spreading properties
Given a measurable function g on $\mathbb{R}^{3}$, suppose that the collision operator satisfies Assumption B with $\nu<0$. Then, there exists $C_{g}^{L}>0$ which depends only on $n_{b}$$C_{\Phi}$, $\rho_{g}$, and $e_{g}$ (and $l_{g,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(3.1) $|L[g](v)|\leq C_{g}^{L}\langle v\rangle^{\gamma^{+}}.$
@@
We define $C_{L}:=\sup_{(t,x)\in[0,T]\times\mathbb{\mathbb{R}}^{3}}C^{L}_{f(t,x,\cdot)}$.
@@
(3.2) $Q^{+}\left(\mathbf{1}_{B(v_{0},R)},\mathbf{1}_{B(v_{0},r)}\right)\geq Cl_{b}c_{\Phi}R^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\sqrt{r^{2}+R^{2}}(1-\xi))},$
@@
In particular, when $\delta=r=R$, we have
(3.3) $Q^{+}\left(\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right)\geq Cl_{b}c_{\Phi}\delta^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))}.$
@@
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$. Suppose that $f(t,x,v)$ is a continuous mild solution to (1.1)–(1.3) with $M>0$ and $0<E_{f}<\infty$. Then, there exist $x_{I}\in\Omega$ with $\Delta^{0}\in(0,\min\{\frac{d}{10},\frac{1}{2}d(x_{I},\partial\Omega)\}]$, $v_{I}\in\mathbb{R}^{3}$, $\alpha^{\prime}_{0}>0$, which depend on $\Omega,M,E_{f}$ and the modulus of continuity of $f$, such that for all $t\in[0,\Delta^{0}]$, $x\in B(x_{I},\Delta^{0})\cap\Omega$, we have
@@
(3.5) $\int_{\Omega}\int_{\mathbb{R}^{3}}(|x|^{2}+|v|^{2})f(t,x,v)dxdv\leq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f}<\infty.$
@@
(3.6) $\int_{B(0,R_{0})}\int_{B(0,R_{0})\cap\Omega}f_{0}(x,v)dxdv\geq\frac{M}{2}>0.$
@@
(3.7) $\begin{split}&\int\int_{\Omega\times\mathbb{R}^{3}-(\Omega\cap B(0,R_{0}))\times B(0,R_{0})}(|x|^{2}+|v|^{2})f_{0}(x,v)dxdv\\ >&R_{0}^{2}\frac{M}{2}\geq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f},\end{split}$
@@
Hence, we can take a point $(x_{I},v_{I})\in\Omega\times B(0,R_{0})$ such that $f(0,x_{I},v_{I})>\frac{M}{3|B(0,R_{0})||B(0,R_{0})\cap\Omega|}$. We define $\widetilde{\Delta}^{0}:=\frac{d(x_{I},\partial\Omega)}{8R_{0}}$. By continuity of $f(t,x,v)$ on $0,\widetilde{\Delta}^{0}]\times\overline{B(x_{I},\frac{d(x_{I},\partial\Omega)}{4})}\times\overline{B(v_{I},R_{0})}\subset\Gamma_{conti}$, there exists $\Delta^{0}\in(0,\min\{\frac{d}{10},\frac{1}{2}d(x_{I},\partial\Omega)]$ which depends on the modulus of continuity of $f$ such that $B(x_{I},\Delta^{0})\subset\Omega$ and
(3.8) $f(t,x,v)\geq\frac{M}{4|B(0,R_{0})||B(0,R_{0})\cap\Omega|},\ \forall(t,x,v)\in[0,\Delta^{0}]\times B(x_{I},\Delta^{0})\times B(v_{I},\Delta^{0}).$
@@
(3.11) $\begin{split}&\alpha_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)\\ :=&Cr_{n}^{3+\gamma}(\Delta_{2})\int_{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}e^{-(t-s)C<2r_{n}(\Delta_{2})+|v^{\prime}|>^{\gamma^{+}}}\alpha_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}ds\\ =&Cr_{n}^{3+\gamma}(\Delta_{2})\int_{0}^{t-t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}e^{-uC<2r_{n}(\Delta_{2})+|v^{\prime}|>^{\gamma^{+}}}\alpha_{n}(\tau,t-u,\Delta_{2},A,|v^{\prime}|)^{2}du.\end{split}$
For $\gamma\geq 0$, $C$ is a constant which depends on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $C_{\Phi}$ and $E_{f}$. For $\gamma<0$, the constant $C$ also depends on $L_{f,p}$, where $p>\frac{3}{3+\gamma}$.
@@
We refer to Lemma 3.3, [1]. ∎
Remark 7.
@@
Here, the functions $m^{\prime}(\tau,\tau^{\prime},\Delta_{2},B,D)$, $\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,E)$, $\mathbb{V}^{\prime}(\tau,\tau^{\prime},y,x^{\prime})$, and $\mathbb{D}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2})$ are defined as change the notation BCD
$m^{\prime}(\tau,\tau^{\prime},\Delta_{2},B,D):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+B+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{D}}\right\rceil\right\},$
$\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,E):=\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,E\right)e^{-C_{L}\left<\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right>^{\gamma^{+}}},$
@@
Here, $C_{L}$ is defined in Definition 3.1, $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,E\right)$ is as defined in Proposition 3.2, and we denote $d_{\Omega}:=\text{\rm diam}(\Omega)$ as the diameter of $\Omega$.
@@
Now, we prove the lower bound (3.12).
As long as $x\in B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)$ and $v\in B\left(v^{\prime\prime},\min\left\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\right\}\right)$, for

### v4

@@ 2. Geometric properties near the boundary
Furthermore, we have $|\nabla\phi_{i}|<\frac{1}{100}$, $|\nabla^{2}\phi_{i}|<\widetilde{C}(\Omega)$ for some $\widetilde{C}(\Omega)>\max\left\{\frac{4}{d},\frac{1}{2d_{r}}\right\}$, and for any $1\leq i\leq m_{1}$,
@@
Remark 2.1.
It is known that any bounded domain with $C^{2}$ boundary satisfies the uniform interior sphere condition. That is, there exists $d_{r}$ such that for any $x\in\partial\Omega$, one can find a ball with radius $d_{r}>0$ in $\Omega$ that intersects with $\partial\Omega$ only on $x$. In this article, we choose $d$ from Lemma 2.1 to be smaller than $d_{r}$.
@@
This is a direct consequence of the fact that $|\nabla\phi_{i}|<\frac{1}{100}$ and the mean value theorem. Indeed, given $x=x_{i}^{0}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x_{i}^{0})$, if there exists $x^{\prime}=x_{i}^{0}+x^{\prime}_{1}e^{1}_{i}+x^{\prime}_{2}e^{2}_{i}-x^{\prime}_{3}n(x_{i}^{0})\in B(x,\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2})))$ such that $x^{\prime}\in\partial\Omega\cap B(x^{0}_{i},3d))$, we have by Lemma 2.1 that
@@
Remark 2.2.
Notice that by the triangle inequality we have $\Omega_{\frac{d}{8}}\subset\bigcup\limits_{i=1}^{m_{1}}B(x_{i}^{0},\frac{d}{4})$, where $\Omega_{\epsilon}:=\{x\in\Omega|\ d(x,\partial\Omega)<\epsilon\}$.
@@
By Lemma 2.1 and Remark 2.2, there exist $m_{1}$ and $\{x_{i}^{0}\}_{i=1}^{m_{1}}$ such that
@@
Clearly, we have $z^{\prime\prime}-z_{q_{1}}=W_{1,//}+W_{1,\perp}$, $W_{1,//}\mathrel{\!/\mkern-5.0mu/\!}Z_{2}\mathrel{\!/\mkern-5.0mu/\!}x-y$, and $W_{1,\perp}\perp Z_{2}$.
@@
Suppose that $\Omega$ satisfies Assumption A with $\delta$, $d$ as given in Lemma 2.1. Then, $\Omega-\Omega_{\frac{d}{8}}$ is pathwise connected; that is, given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$, there exists a continuous curve in $\Omega-\Omega_{\frac{d}{8}}$ joining $x^{\prime}_{1}$ and $x^{\prime}_{2}$.
@@
Given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$. Since the domain $\Omega$ is connected, $\Omega$ is pathwise connected. Hence, there exist $l\in\mathbb{N},\{\widetilde{x}_{i}\}_{i=1}^{l}\in\Omega$ such that the following curve
@@
As Remark 2.2 shows, we have
@@
We also notice that, using the same argument as (2.11), we have $|n(x^{0}_{k(j)})-n(x^{0}_{k(j+1)})|<\frac{1}{50}$ for $1\leq j\leq l-1$, from which we deduce that
@@
all the assumptions of Lemma 2.3 are satisfied. Hence, we deduce that
@@
for $i=1,2$. Indeed, if $x^{\prime}_{i}\in\Omega_{\frac{d}{8}}$, then by Lemma 2.1, there exists $1\leq l(i)\leq m_{1}$ such that $x^{\prime}_{i}\in B(x_{l(i)}^{0},\frac{d}{4})$ and we can parameterize the boundary of $\Omega$ near $x^{0}_{l(i)}$ as follows:
@@
By the mean value theorem, there exists $\zeta^{\prime}$ such that
@@
(2.20) $\begin{split}&\mathbb{O}(w_{1}):=\bigg\{y\in\mathbb{Y}\mid\mathrm{There\ exists\ a\ zigzag}\,\ \{w_{1},u_{1},...,u_{s},y\}\ni\,u_{i}\in\mathbb{Y}\,\forall\,i=1,...,s,\\ &d(\overline{w_{1}u_{1}},\partial\Omega)>\frac{d}{10},d(\overline{u_{i}u_{i+1}},\partial\Omega)>\frac{d}{10}\forall\,1\leq i\leq s-1,d(\overline{u_{s}y},\partial\Omega)\>\frac{d}{10}\bigg\}.\end{split}$
@@
$\overline{\xi}:=\sup\left\{0\leq s\leq 1\mid d(\mathbb{O}(w_{1}),\overline{\phi}(s))\leq\frac{d}{15}\right\}.$
@@
Thus, we have that $w_{4}\in\mathbb{O}(w_{1})$. This implies that $\frac{d}{16}\leq d(\mathbb{O}(w_{1}),w_{3})\leq d(w_{4},w_{3})<\frac{d}{16}$, which is a contradiction. We conclude that $\Omega-\Omega_{\frac{d}{8}}\subset B\left(\mathbb{O}(w_{1}),\frac{d}{16}\right)$. Therefore, $\mathbb{O}(w_{1})$ contains $\mathbb{Y}$, and $w_{2}\in\mathbb{O}(w_{1})$, which shows that there exists a zigzag between $w_{1}$ and $w_{2}$ satisfying the conditions mentioned in Lemma 2.5. ∎
@@
We invoke the Remark 2.1 and use the uniform interior sphere condition to deduce that for any $0\leq s\leq\frac{\sin{\theta_{i}}}{\widetilde{C}(\Omega)}$, we have
@@
Here, we use the fact that $|v|\leq v_{M}$. For the (2.30), we again use the uniform interior sphere condition to deduce that for any $y\in\partial\Omega$ and $w\in\mathbb{R}^{3}$ with $n(y)\cdot w<0$, $\theta_{w}:=\frac{\pi}{2}-\arccos{|n(y)\cdot w|}$, we have
@@ 3. Operator estimates and spreading properties
Given a measurable function g on $\mathbb{R}^{3}$, suppose that the collision operator satisfies Assumption B with $\nu<0$. Then, there exists $C_{g}^{L}>0$ which depends only on $n_{b}$, $C_{\Phi}$, $\rho_{g}$, and $e_{g}$ (and $l_{g,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(3.1) $|L[g](v)|\leq C_{g}^{L}\left\langle v\right\rangle^{\gamma^{+}}.$
@@
We define $C_{L}:=\sup_{(t,x)\in[0,T]\times\mathbb{R}^{3}}C^{L}_{f(t,x,\cdot)}$.
@@
(3.2) $Q^{+}\left[\mathbf{1}_{B(v_{0},R)},\mathbf{1}_{B(v_{0},r)}\right]\geq Cl_{b}c_{\Phi}R^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\sqrt{r^{2}+R^{2}}(1-\xi))},$
@@
In particular, when $r=R=\delta$, we have
(3.3) $Q^{+}\left[\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right]\geq Cl_{b}c_{\Phi}\delta^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))}.$
@@
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$. Suppose that $f(t,x,v)$ is a continuous mild solution to (1.1)–(1.3) with $M>0$ and $0<E_{f}<\infty$. Then, there exist $x_{I}\in\Omega$ with $\Delta^{0}\in\left(0,\;\min\left\{\frac{d}{10},\,\frac{1}{2}d(x_{I},\partial\Omega)\right\}\right]$, $v_{I}\in\mathbb{R}^{3}$, $\alpha^{\prime}_{0}>0$, which depend on $\Omega,M,E_{f}$ and the modulus of continuity of $f_{0}$, such that for all $t\in[0,\Delta^{0}]$, $x\in B(x_{I},\Delta^{0})\cap\Omega$, we have
@@
(3.5) $\int_{\mathbb{R}^{3}}\int_{\Omega}(|x|^{2}+|v|^{2})f(t,x,v)\,dx\,dv\leq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f}<\infty.$
@@
(3.6) $\int_{B(0,R_{0})}\int_{B(0,R_{0})\cap\Omega}f_{0}(x,v)\,dx\,dv=\int_{B(0,R_{0})}\int_{\Omega}f_{0}(x,v)\,dx\,dv\geq\frac{M}{2}>0.$
@@
(3.7) $\begin{split}&\int\int_{\Omega\times\mathbb{R}^{3}-(B(0,R_{0})\cap\Omega)\times B(0,R_{0})}(|x|^{2}+|v|^{2})f_{0}(x,v)\,dx\,dv\\ >&R_{0}^{2}\frac{M}{2}\geq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f},\end{split}$
@@
Hence, we can take a point $(x_{I},v_{I})\in\Omega\times B(0,R_{0})$ such that $f(0,x_{I},v_{I})>\frac{M}{3|B(0,R_{0})||\Omega|}$. By the uniform continuity of $f_{0}(x,v)$ on $\overline{\Omega}\times\overline{B(0,2R_{0})}$, there exist $0<\Delta_{I}<\min\left\{\frac{d(x_{I},\partial\Omega)}{2},R_{0}\right\}$ such that
$f_{0}(x,v)\geq\frac{M}{6|B(0,R_{0})||\Omega|}$
for any $(x,v)\in B(x_{I},\Delta_{I})\times B(v_{I},\Delta_{I})$.
Next, we observe that for any $(t,x,v)\in\left[0,\frac{\Delta_{I}}{4R_{0}}\right]\times B\left(x_{I},\frac{\Delta_{I}}{2}\right)\times B(v_{I},\Delta_{I})$, we have
$(x-tv,v)\in B(x_{I},\Delta_{I})\times B(v_{I},\Delta_{I})$
. As a result, we deduce from (1.12) and Lemma 3.1 that
(3.8) $\begin{split}f(t,x,v)&\geq f_{0}(X_{0,t}(x,v),v)\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &\geq f_{0}(x-tv,v)e^{-tC_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ &\geq\frac{M}{6|B(0,R_{0})||\Omega|}e^{-\frac{\Delta_{I}C_{L}}{4R_{0}}\left\langle 2R_{0}\right\rangle^{\gamma^{+}}}\end{split}$
for any $(t,x,v)\in\left[0,\frac{\Delta_{I}}{4R_{0}}\right]\times B\left(x_{I},\frac{\Delta_{I}}{2}\right)\times B(v_{I},\Delta_{I})$, where the constant $C_{L}$ depends only on $n_{b}$$C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
We conclude the proof by defining
$\Delta^{0}:=\min\left\{\frac{\Delta_{I}}{4R_{0}},\frac{\Delta_{I}}{2},\frac{d}{10}\right\}.$
@@
(3.11) $\begin{split}&\alpha_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)\\ :=&\frac{C_{Q}}{2}r_{n}^{3+\gamma}(\Delta_{2})\int_{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}e^{-(t-s)C_{L}\left\langle 2r_{n}(\Delta_{2})+|v^{\prime}|\right\rangle^{\gamma^{+}}}\alpha_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}ds\\ =&\frac{C_{Q}}{2}r_{n}^{3+\gamma}(\Delta_{2})\int_{0}^{t-t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}e^{-uC_{L}\left\langle 2r_{n}(\Delta_{2})+|v^{\prime}|\right\rangle^{\gamma^{+}}}\alpha_{n}(\tau,t-u,\Delta_{2},A,|v^{\prime}|)^{2}du.\end{split}$
@@
The following proof is also a summary of the proof of Lemma 3.3 in [1]. We prove the proposition by induction on $n$. The case $n=0$ is exactly the assumption. Assume that Proposition 3.2 holds for $n=k$. Given $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{k+1}})$, $v\in B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$, we first notice that when $s\in\left[\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(|v^{\prime}|+2r_{k}(\Delta_{2}))}\right\},t\right]$, we have
which implies that $X_{s,t}(x,v)\in B(x^{\prime},\frac{\Delta_{2}}{2^{k}})\subset\Omega$.
We consider the second term of the right hand side of (1.12) to obtain the following lower bound:
(3.12) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v))\,ds^{\prime}\right)Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Furthermore, by Lemma 3.1, we have
(3.13) $|L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)|\leq C_{L}\langle v\rangle^{\gamma^{+}}\leq C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}},$
where the constant $C_{L}$ depends only on $n_{b}$$C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
Hence, the following estimate:
(3.14) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Since we have $|X_{s,t}(x,v)-x^{\prime}|\leq\frac{\Delta_{2}}{2^{k}}$, we can use the induction hypothesis:
(3.15) $f(s,X_{s,t}(x,v),w)\geq\alpha_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)1_{B(v^{\prime},r_{k}(\Delta_{2}))}(w),\ \forall w\in\mathbb{R}^{3}$
to deduce that
(3.16) $\begin{split}&f(t,x,v)\\ \geq&\int_{\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}(\Delta_{2})+|v^{\prime}|)}\right\}}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)Q^{+}[1_{B(v^{\prime},r_{k}(\Delta_{2}))}(\cdot),1_{B(v^{\prime},r_{k}(\Delta_{2}))}(\cdot)](v)\,ds.\end{split}$
Then we use Lemma 3.2 to ”spread” the lower bound:
For any $\xi\in(0,1)$ and $w\in\mathbb{R}^{3}$
(3.17) $Q^{+}\left[1_{B(v^{\prime},r_{k}(\Delta_{2}))},1_{B(v^{\prime},r_{k}(\Delta_{2}))}\right](w)\geq C_{Q}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}1_{B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))}(w).$
Hence, for any $\xi\in(0,1)$
(3.18) $\begin{split}&f(t,x,v)\\ \geq&\int_{\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}(\Delta_{2})+|v^{\prime}|)}\right\}}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)C_{Q}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}1_{B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))}(v)\,ds.\end{split}$
Set $\xi=\frac{1}{4}$ and notice that $B(v^{\prime},r_{k+1}(\Delta_{2})))\subset B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$. We conclude thus the proof of Proposition 3.2.
Remark 3.2.
@@
Here, the functions $m^{\prime}(\tau,\tau^{\prime},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$, $\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3})$, $\mathbb{V}^{\prime}(\tau,\tau^{\prime},y,x^{\prime})$, and $\mathbb{D}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2})$ are defined as
$m^{\prime}(\tau,\tau^{\prime},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2}):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+\mathfrak{a}_{1}+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{\mathfrak{y}_{2}}}\right\rceil\right\},$
$\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3}):=\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,\mathfrak{y}_{3}\right)e^{-C_{L}\left\langle\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right\rangle^{\gamma^{+}}},$
@@
Here, $C_{L}$ is defined in Definition 3.1, $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,\mathfrak{y}_{3}\right)$ is as defined in Proposition 3.2, and we denote $d_{\Omega}:=\text{\rm diam}(\Omega)$ as the diameter of $\Omega$.
@@
Now, we prove the lower bound (3.19). As long as $x\in B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)$ and $v\in B\left(v^{\prime\prime},\min\left\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\right\}\right)$, for